At the annual christmas party a drawing was held for: 20 gift certificates worth $25 each, 40 turkeys, and 40 boxes of candy.There are 50 employees so that each has two draws.
a) what is the probability the first employee drew 2 gift certificates?
b) what is the probability he drew a gift certificate and then a certificate for a turkey?
c) what is the probability he drew a certificate for a turkey and then for a box of candy?

Answer by Boreal(15235)

2 gift certificates in a row is 20/100*19/99=(1/5)(19/99)=19/495.
gift certificate and turkey is 20/100*40/99=(1/5)(40/99)=40/495=8/99
turkey and candy is 40/100*40/99=(2/5)*(40/99)=80/495=16/99